# Sharp Online Hardness for Large Balanced Independent Sets

### v1

@@ 1.4 Open Problems
Future Queries
This naturally raises the following question: can algorithms with limited future information outperform the $(1+\epsilon)\alpha_{\rm COMP}$ threshold for $G_{\mathrm{bip}}(n,p)$? We conjecture that the answer is yes:
Conjecture 1.7.
Let $E$ be the set of all future edges ever revealed to the algorithm. For any $\epsilon$, there exists a constant $c_{\epsilon}>0$ and an online algorithm $\mathcal{A}$ that finds a $\gamma$-balanced independent set of size at least $(1+\epsilon)\alpha_{\rm COMP}$ whp, provided
$\left|E\cap\binom{\mathcal{A}(G)}{2}\right|\leq c_{\epsilon}\log_{b}^{2}(np).$
Remark 1.8.
Importantly, an online algorithm resolving Conjecture 1.7 cannot be fully oblivious to the vertex arrival order. Suppose that it produces a $\gamma$-balanced independent set of size $(1+\epsilon)\alpha_{\rm COMP}$, and the first $n$ arrivals all come from the same side of the bipartition. By the online nature, decisions cannot be deferred: at least $(1+\epsilon)\log_{b}n$ vertices from that side must be included. This, however, would prevent adding any vertices from the opposite side as the expected number of vertices there with no edge to the chosen $(1+\epsilon)\log_{b}n$ vertices is only $n^{-\epsilon}$.
@@ 4. Case 4:
7:   if $|I_{t-1}\cap R_{t}|<\frac{(1-\gamma)}{\gamma}\,(1-\epsilon)\log_{b}n$ and $N_{t}(v_{t})\cap I_{t-1}=\emptyset$ then
@@
$\mathbb{P}[N_{t}(v_{t})\cap I_{t-1}=\emptyset]$ $=\mathbb{P}\left[\mbox{\sf Bin}(|I_{t-1}\cap R|,p)=0\mid v_{t}\in L\right]+\mathbb{P}\left[\mbox{\sf Bin}(|I_{t-1}\cap L|,p)=0\mid v_{t}\in R\right]$
@@
  • Initialize $\tilde{I}_{0}=\emptyset$.
@@
Let us now analyze Algorithm 5 conditioning on the event $T_{f}\leq n^{1-\epsilon/2}$. For brevity, we avoid stating this conditioning at every step of the proof. Once again, we will consider an alternate process $\bar{I}$. Note that during a single iteration of the for loop of Algorithm 5, we do nothing if $v_{t}\in L$ or $|I_{t-1}\cap R|=\frac{(1-\gamma)}{\gamma}(1-\epsilon)\log_{b}n$. Suppose the latter condition first occurs at timestep $T_{b}$. From this time point, we continue the algorithm as follows. Let $\bar{I}_{T_{b}-1}\coloneqq I_{T_{b}-1}$. At each time step $t\geq T_{b}$, we do the following: if $v_{t}\in L$, we do nothing; if $v_{t}\in R$, we add $v_{t}$ to $\bar{I}_{t-1}$ if $N_{t}(v_{t})\cap\bar{I}_{t-1}=\emptyset$.
@@ 2. Case 2:
pages10 rangepages10 rangepages10 rangepages10 rangepages-1 rangepages9 rangepages6 rangepages1 rangepages32 rangepages14 rangepages9 rangepages56 rangepages10 rangepages30 rangepages5 rangepages10 rangepages26 rangepages10 rangepages33 rangepages67 rangepages12 rangepages31 rangepages51 rangepages1 rangepages8 rangepages24 rangepages30 rangepages12 rangepages38 rangepages28 rangepages169 rangepages13 rangepages10 rangepages11 rangepages14 rangepages1 rangepages1 rangepages22 rangepages35 rangepages27 rangepages36 rangepages31 rangepages-1

### v2

@@ 1.2 Summary of Main Results
Future Queries
This naturally raises the following question: can algorithms with limited future information outperform the $(1+\epsilon)\alpha_{\rm COMP}$ threshold for $G_{\mathrm{bip}}(n,p)$? We show that the answer is yes:
Theorem 1.7 (Informal, see Theorem 3.8).
For any $\epsilon>0$, there exists an online algorithm which makes limited future queries and finds a $\gamma$-balanced independent set of size $(1+\epsilon)\alpha_{\rm COMP}$ with probability at least $1-\exp(-n^{\Theta(1)})$.
Our algorithm is online, and runs in super-polynomial time. It has a mild dependence on the vertex arrival order—indeed, no online algorithm that is fully oblivious to arrival order can surpass $(1+\epsilon)\alpha_{\rm COMP}$. See Section 3.3 for further discussion.
@@ 3.2 Algorithmic Results
3.3 Surpassing $\alpha_{\rm COMP}$ with Limited Future Queries
Our algorithmic lower bound shows that online algorithms, operating exclusively based on the information available up to round $t$, cannot surpass $\alpha_{\rm COMP}$. This provides strong evidence for the conjecture that $\alpha_{\rm COMP}$ is the true computational threshold for this model.
At the same time, prior work [GKW25] made an intriguing observation. For the largest independent set problem in $G(n,\tfrac{1}{2})$, they showed that granting the algorithm access to a limited amount of additional information allows it to exceed the computational threshold. [^5] The resulting algorithm, albeit being online, requires super-polynomial time. This naturally raises the following question: for the balanced independent set problem in $G_{\mathrm{bip}}(n,p)$, can $\alpha_{\rm COMP}$ be surpassed if the algorithm is permitted a limited number of future queries at each round?
In this section, we show that the answer is yes: online algorithms augmented with a small number of future queries can indeed surpass $\alpha_{\rm COMP}$. To make this precise, we extend the standard definition of an online algorithm (Definition 3.2) to allow future queries
Definition 3.7.
For $c>0$, let $\mathcal{C}_{c}$ be the class of online algorithms operating according to Definition 3.2, with the following extension. At each round $t\in[2n]$, the algorithm may, in addition to observing the edges $(v_{t},v)$ for $v\in L_{t}\cup R_{t}$, query a (possibly random) set $S_{t}$ of vertex pairs $\{i,j\}$—which may include vertices not yet revealed—and reveal the status of all edges in $S_{t}$. The decision is then based on the combined information, where the number of future queries satisfy
$\Biggl{|}\;\bigcup_{t=1}^{2n}S_{t}\;\cap\;\binom{\mathcal{A}(G_{\mathrm{bip}}(n,p))}{2}\Biggr{|}\;\leq\;c(\log_{b}n)^{2}.$ (3)
Note that $S_{t}$ may include pairs involving vertices from $v_{t+1},\dots,v_{2n}$}. For this reason, we refer to the edges in $S_{t}$ as future edges. The total amount of such future information is limited to $O(\log^{2}n)$.
Our final main result is as follows.
Theorem 3.8.
For any $\epsilon>0$ and $p,\gamma\in(0,1)$, there exists $c\coloneqq c(\gamma,\epsilon)$ and an online algorithm $\mathcal{A}\in\mathcal{C}_{c}$ that $(k,\delta)$-optimizes the $\gamma$-balanced independent set problem in $G_{\mathrm{bip}}(n,p)$, where
$k=(1+\epsilon)\alpha_{\rm COMP},\qquad\text{and}\qquad\delta=1-\exp(-n^{\Theta(1)}).$
Our algorithm proceeds in three phases, described informally as follows. The first phase runs for $T=o(n)$ rounds and greedily constructs a $\gamma$-balanced independent set $I_{T}=R_{T}\cup L_{T}$ with $R_{T}\subset R$ and $L_{T}\subset L$, of size $c^{\prime}\log_{b}n/\gamma$ for a suitable constant $c^{\prime}\coloneqq c^{\prime}(\gamma,\epsilon)$. The second phase is an exploration phase, where—using future queries—it identifies sets $W_{L}\subset L$ and $W_{R}\subset R$ not yet inspected, such that there are no edges (i) between $R_{T}$ and $W_{L}$, and (ii) between $L_{T}$ and $W_{R}$. In the third phase, it performs a brute-force search to identify a $\gamma$-balanced independent set of size $(1+\epsilon-c^{\prime})\log_{b}n/\gamma$ inside $G_{\mathrm{bip}}(W_{L}\cup W_{R},p)$, and augments this to $I_{T}$.
Our analysis shows that the procedure articulated above can be implemented in an online fashion; see Section 7 for the details.
Importantly, our algorithm is not fully oblivious to the vertex arrival order. In fact, no online algorithm can produce a $\gamma$-balanced independent set of size $(1+\epsilon)\log_{b}n$ (whp) while remaining completely oblivious to the arrival order. Suppose, for contradiction, that such an algorithm outputs a $\gamma$-balanced independent set of size $(1+\epsilon)\alpha_{\rm COMP}$. If the first $n$ arrivals all lie on the same side of the bipartition, then by the online constraint (decisions cannot be deferred), the algorithm must select at least $(1+\epsilon)\log_{b}n$ vertices from that side. This, however, precludes adding vertices from the opposite side, since the expected number of vertices there with no edge to the chosen $(1+\epsilon)\log_{b}n$ vertices is at most $n^{-\epsilon}$. Thus, some dependence on arrival order is unavoidable. That said, our algorithm’s dependence on arrival order is fairly mild: the only step where order matters is in the first greedy phase (Algorithm 7), designed specifically to avoid such pathological situations.
Number of Future Queries
Our analysis in Section 7 shows that it suffices to take
$c(\gamma,\epsilon)\coloneqq\frac{1-\gamma}{\gamma}\left((1+\epsilon)^{2}-(c^{\prime})^{2}\right),$
where $c^{\prime}$ is any arbitrary constant satisfying
$c^{\prime}<\frac{\gamma\bigl{(}1-(1+\epsilon)(1-\gamma)\bigr{)}}{(1-\gamma)^{2}}.$
In the balanced case $(\gamma=\tfrac{1}{2}$), the condition on $c^{\prime}$ boils down to $c^{\prime}<1-\epsilon$. In this case, one can choose $c(\frac{1}{2},\epsilon)$ such that $c(\frac{1}{2},\epsilon)\to 0$ as $\epsilon\to 0$, e.g., by taking $c(\frac{1}{2},\epsilon)=6\epsilon$. By contrast, for $\gamma<\tfrac{1}{2}$, our analysis suggests that it is not possible to choose $c(\gamma,\epsilon)$ such that $c(\gamma,\epsilon)\to 0$ as $\epsilon\to 0$. At first glance, this may seem surprising. However, note the following: as $\gamma\to\frac{\epsilon}{1+\epsilon}$, we have $\alpha_{\rm COMP}\to\alpha_{\rm STAT}$. In particular, as the statistical-computational gap is “smaller” for smaller values of $\gamma$, one expects surpassing the gap to be somewhat “harder”. Indeed, our analysis reflects this: interpolating between $\gamma=1/2$ and $\gamma=\frac{\epsilon}{1+\epsilon}$, we shift from $c(\gamma,\epsilon)=\Theta(\epsilon)$ to $c(\gamma,\epsilon)=\Theta(1/\epsilon)$. [^6] We remark that the distinction between $\gamma=1/2$ and $\gamma<1/2$ is prevalent in the analogous setting of $\gamma$-balanced colorings. In fact, while the problem has been heavily studied for $\gamma=1/2$ [Cha23, FK10, Dha25, DW25], no results are known for $\gamma$-balanced colorings for $\gamma<1/2$ (see the discussion in [DW25, Section 1.3] on the challenges involved).
@@ 4. Case 4:
7:   if $|I_{t-1}\cap R_{t}|<\frac{(1-\gamma)}{\gamma}\,(1-\epsilon)\log_{b}n$ and $N_{t}(v_{t})\cap I_{t-1}=\varnothing$ then
@@
$\mathbb{P}[N_{t}(v_{t})\cap I_{t-1}=\varnothing]$ $=\mathbb{P}\left[\mbox{\sf Bin}(|I_{t-1}\cap R|,p)=0\mid v_{t}\in L\right]+\mathbb{P}\left[\mbox{\sf Bin}(|I_{t-1}\cap L|,p)=0\mid v_{t}\in R\right]$
@@
  • Initialize $\tilde{I}_{0}=\varnothing$.
@@
Let us now analyze Algorithm 5 conditioning on the event $T_{f}\leq n^{1-\epsilon/2}$. For brevity, we avoid stating this conditioning at every step of the proof. Once again, we will consider an alternate process $\bar{I}$. Note that during a single iteration of the for loop of Algorithm 5, we do nothing if $v_{t}\in L$ or $|I_{t-1}\cap R|=\frac{(1-\gamma)}{\gamma}(1-\epsilon)\log_{b}n$. Suppose the latter condition first occurs at timestep $T_{b}$. From this time point, we continue the algorithm as follows. Let $\bar{I}_{T_{b}-1}\coloneqq I_{T_{b}-1}$. At each time step $t\geq T_{b}$, we do the following: if $v_{t}\in L$, we do nothing; if $v_{t}\in R$, we add $v_{t}$ to $\bar{I}_{t-1}$ if $N_{t}(v_{t})\cap\bar{I}_{t-1}=\varnothing$.
@@ 2. Case 2:
7 Future Queries: Proof of Theorem 3.8
Fix an arbitrary constant
$c^{\prime}<\frac{\gamma\bigl{(}1-(1+\epsilon)(1-\gamma)\bigr{)}}{(1-\gamma)^{2}},$ (22)
where $\epsilon<\frac{\gamma}{1-\gamma}$. Indeed, otherwise there exists no $\gamma$-balanced independent set of size $(1+\epsilon)\alpha_{\rm COMP}$ per Theorem 3.1. Note that as $(1+\epsilon)(1-\gamma)\geq 1-\gamma$ and $\gamma\leq\frac{1}{2}$, we have
$c^{\prime}<\frac{\gamma^{2}}{(1-\gamma)^{2}}\leq 1.$ (23)
In what follows, we prove the existence of such an online algorithm with
$c(\gamma,\epsilon)=\frac{1-\gamma}{\gamma}\left((1+\epsilon)^{2}-(c^{\prime})^{2}\right).$
As noted, our algorithm proceeds in three phases.
Phase I: Greedy Rounds
In phase I, we find a $\gamma$-balanced independent set of size $c^{\prime}\log_{b}n/\gamma$ for $c^{\prime}$ arising in (22). Proceeding analogously to the proof of Theorem 3.4, we first run Algorithm 5 until time $T_{1}$ where
$T_{1}\coloneqq\min\bigl{\{}t\geq 1:\max\{|I_{t}\cap L|,|I_{t}\cap R|\}=c^{\prime}\log_{b}n\bigr{\}}.$
Plugging $1-c^{\prime}$ in place of $\epsilon$ in Lemma 5.1, we deduce
$\mathbb{P}\left[T_{1}\leq n^{1-(1-c^{\prime})/2}\right]=1-\exp\left(-\Omega\left(n^{(1-c^{\prime})/2}\right)\right).$ (24)
The bound (23) ensures that (24) is not vacuous. Note that at time $T_{1}$, one of the partitions to the independent set $I_{T_{1}}$ is of size $c^{\prime}\log_{b}n$. By relabeling the partitions if necessary, we again assume this partition is $L$. For the second stage, we employ the following modification of Algorithm 5. Set $T_{2}=n^{c^{\prime}+\delta}$ where $0<\delta<1-c^{\prime}$ is arbitrary.
Algorithm 3 Phase I: Part II
1:for $t=T_{1}+1,\ldots,T_{1}+T_{2}$ do
2:  Sample a random vertex $v_{t}\in R\setminus R_{t-1}$.
3:  $L_{t}=L_{t-1}$, $R_{t}=R_{t-1}\cup\{v_{t}\}$.
4:  if $|I_{t-1}\cap R|<c^{\prime}\log_{b}n$ and $N_{t}(v_{t})\cap I_{t-1}=\varnothing$ then
5:   Set $I_{t}=I_{t-1}\cup\{v_{t}\}$.
6:  else
7:   $I_{t}=I_{t-1}$.
8:  end if
9:end for
Lemma 7.1.
$\mathbb{P}\left[\left|I_{T_{1}+T_{2}}\cap R\right|=c^{\prime}\log_{b}n\right]=1-\exp\left(-\Omega(n^{\delta})\right)$.
Proof.
For $\{v_{T_{1}+1},\dots,v_{T_{1}+T_{2}}\}\subset R$, let $N$ denote the number of $i$ such that $T_{1}+1\leq i\leq T_{1}+T_{2}$ for which there are no edges between $v_{i}$ and $I_{T_{1}}$. We have
$N=\sum_{i=T_{1}+1}^{T_{1}+T_{2}}I_{i}\sim{\rm Bin}(T_{2},n^{-c^{\prime}})$
as $I_{i}$ are iid Bernouli random variables with parameter $n^{-c^{\prime}}$. Clearly, it suffices to show that
$\mathbb{P}\bigl{[}N\geq c^{\prime}\log_{b}n\bigr{]}=1-\exp\left(-\Omega\left(n^{\delta}\right)\right),$
as we only add vertices as long as $|I_{t-1}\cap R|<c^{\prime}\log_{b}n$. As $\mathbb{E}[N]=n^{\delta}$, we have
$\mathbb{P}\bigl{[}N<c^{\prime}\log_{b}n\bigr{]}\leq\mathbb{P}\left[N\leq\mathbb{E}[N]/2\right]\leq e^{-\mathbb{E}[N]/8}=e^{-n^{\delta}/8}$
by a standard Chernoff bound (see, e.g., [AS16]), completing the proof. ∎
Combining (24) and Lemma 7.1 through a union bound, we have the following with probability $1-\exp\left(-\Omega(n^{\min\{\delta,(1-c^{\prime})/2\}})\right)$ for $T=T_{1}+T_{2}$: (i) $T=o(n)$, (ii) $I_{T}=I^{(L)}_{T}\cup I^{(R)}_{T}$ is a $\gamma$-balanced independent set of size $c^{\prime}\log_{b}n/\gamma$ where
[…]
Namely, a $\gamma$-balanced independent set of size $c^{\prime}\log_{b}n/\gamma$ is found in $o(n)$ time with high probability. This concludes Phase I of our algorithm. Note that for $t\leq T=T_{1}+T_{2}$, $S_{t}=\varnothing$—i.e., no future queries have been made.
In what follows, we condition on the outcome of the greedy rounds.
Phase II: Exploration Phase
We condition on the outcome of Phase I and fix a $\theta\in(0,1)$ with
$(1-\gamma)(1+\epsilon-c^{\prime})<\theta<1-c^{\prime}\frac{1-\gamma}{\gamma}.$ (25)
Observe that $\theta$ is well defined if
$1-c^{\prime}\frac{1-\gamma}{\gamma}>(1+\epsilon-c^{\prime})(1-\gamma)$
which is equivalent to (22). At round $T+1$, we do the following:
  • Sample a random vertex $v_{T+1}\in[2n]\setminus(L_{T}\cup R_{T})$. If $v_{T+1}\in R$, update $R_{T+1}=R_{T}\cup\{v_{T+1}\}$ and $L_{T+1}=L_{T}$; otherwise update $R_{T+1}=R_{T}$ and $L_{T+1}=L_{T}\cup\{v_{T+1}\}$.
  • Query $S_{T+1}$ involving vertices $\{i,j\}$ with $i\in I^{(R)}_{T}$ and $j\in L\setminus L_{T}$ and reveal the edge status of all pairs $(i,j)$.
  • Using the edges in the step above, identify a $\bar{W}_{L}\subset L\setminus L_{T}$ of size $|\bar{W}_{L}|=n^{\theta}$ for which $E[\bar{W}_{L},I_{T}^{(R)}]=\varnothing$. Lemma 7.2 below justifies the existence of $\bar{W}_{L}$ whp.
At the end of round $T+1$, we set $I_{T+1}\coloneqq I_{T}$. Similarly, at round $T+2$, we do the following:
  • Sample a random vertex $v_{T+2}\in[2n]\setminus(L_{T+1}\cup R_{T+1})$. If $v_{T+2}\in R$, update $R_{T+2}=R_{T+1}\cup\{v_{T+2}\}$ and $L_{T+2}=L_{T+1}$; otherwise update $R_{T+2}=R_{T+1}$ and $L_{T+2}=L_{T+1}\cup\{v_{T+2}\}$.
  • Query $S_{T+2}$ involving vertices $\{i,j\}$ with $i\in I^{(L)}_{T}$ and $j\in R\setminus R_{T+1}$ and reveal the edge status of all pairs $(i,j)$.
  • Using the edges in the step above, identify a $\bar{W}_{R}\subset R\setminus R_{T+1}$ of size $|\bar{W}_{R}|=n^{\theta}$ for which $E[I_{T}^{(L)},\bar{W}_{R}]=\varnothing$. Lemma 7.2 justifies the existence of $\bar{W}_{R}$ whp.
We then set $I_{T+2}\coloneqq I_{T+1}$.
Lemma 7.2.
For $\theta$ satisfying (25), such sets $\bar{W}_{L}$ and $\bar{W}_{R}$ indeed exist with probability at least $1-\exp\left(-n^{\Theta(1)}\right)$.
Proof.
Let
$N_{L}=\sum_{v\in L\setminus L_{T}}{\mathbb{I}\left\{v\text{ has no neighbors in }I_{T}^{(R)}\right\}}$
Note that for $v\in L\setminus L_{T}$
${\mathbb{I}\left\{v\text{ has no neighbors in }I_{T}^{(R)}\right\}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}{\rm Ber}\left(n^{1-c^{\prime}\frac{1-\gamma}{\gamma}}\right).$
Since we deterministically have $|L_{T}|,|R_{T}|\leq T$, and $T=o(n)$ by conditioning, $N_{L}$ stochastically dominates ${\rm Bin}(n-T,n^{-c^{\prime}(1-\gamma)/\gamma})$, which has mean $n^{1-c^{\prime}(1-\gamma)/\gamma}(1+o(1))$ (see, e.g., [Roc24, Chapter 4]). Proceeding identically to the proof of Lemma 7.1,
$\mathbb{P}\left[N_{L}\geq n^{\theta}\right]\geq\mathbb{P}\left[{\rm Bin}\left(n-T,n^{-c^{\prime}\frac{1-\gamma}{\gamma}}\right)\geq n^{\theta}\right]\geq 1-\exp\left(-n^{\Theta(1)}\right)$ (26)
as $\theta$ satisfies (25). Similarly, define
$N_{R}=\sum_{v\in R\setminus R_{T+1}}{\mathbb{I}\left\{v\text{ has no neighbors in }I_{T}^{(L)}\right\}}.$
Clearly, for $v\in R\setminus R_{T+1}$
${\mathbb{I}\left\{v\text{ has no neighbors in }I_{T}^{(L)}\right\}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}{\rm Ber}\left(n^{1-c^{\prime}}\right).$
As $|R\setminus R_{T+1}|=|R|-|R_{T+1}|\geq n-T-1$, and therefore $N_{R}$ stochastically dominates ${\rm Bin}(n-T-1,n^{-c^{\prime}})$ which has mean $n^{1-c^{\prime}}(1+o(1))$. Proceeding identically as above, we obtain that
$\mathbb{P}\left[N_{R}\geq n^{\theta}\right]\geq 1-\exp\left(-n^{\Theta(1)}\right).$ (27)
Noting that such $\bar{W}_{L}$ and $\bar{W}_{R}$ with sizes $n^{\theta}$ exist if and only if $N_{L},N_{R}\geq n^{\theta}$, we conclude the proof by combining (26) and (27) through a union bound. ∎
Phase III: Brute-Force Step
At round $t=T+3$, we:
  • Sample a vertex $v_{T+3}\in[2n]\setminus(L_{T+2}\cup R_{T+2})$. If $v_{T+3}\in R$, update $R_{T+3}=R_{T+2}\cup\{v_{T+3}\}$ and $L_{T+3}=L_{T+2}$; otherwise update $R_{T+3}=R_{T+2}$ and $L_{T+3}=L_{T+2}\cup\{v_{T+3}\}$.
  • (Pre-Processing) Set $W_{L}\coloneqq\bar{W}_{L}\setminus\{v_{T+1},v_{T+2},v_{T+3}\}$ and $W_{R}\coloneqq\bar{W}_{R}\setminus\{v_{T+1},v_{T+2},v_{T+3}\}$. Note that $|W_{L}-W_{R}|\leq 3$. Removing at most three vertices from the larger set if necessary, we may assume that $|W_{L}|=|W_{R}|=n^{\theta}+O(1)$.
  • (Brute-Force Search) Query $S_{T+3}$ involving vertices $\{i,j\}$ where $i\in W_{L}$ and $j\in W_{R}$ and reveal the edge status of all pairs $(i,j)$. Using this information, identify a $\gamma$-balanced independent set $\mathcal{J}=J_{L}\cup J_{R}$ in $G[W_{L}\cup W_{R}]$ $G_{\mathrm{bip}}(n^{\theta}+O(1),p)$) of size $(1+\epsilon-c^{\prime})\log_{b}n/\gamma$, where
    $|J_{L}|$ $=|\mathcal{J}\cap L|=(1+\epsilon-c^{\prime})\log_{b}n,$
    $|J_{R}|$ $=|\mathcal{J}\cap R|=(1+\epsilon-c^{\prime})\frac{1-\gamma}{\gamma}\log_{b}n.$
    Lemma 7.3 justifies the existence of $\mathcal{J}$ whp. We note that this step has quasi-polynomial running time $n^{O(\log n)}$ as mentioned in Section 3.3.
Set $I_{T+3}\coloneqq I_{T}$.
Lemma 7.3.
The bipartite random graph $G[W_{L}\cup W_{R}]\sim G_{\mathrm{bip}}(n^{\theta}+O(1),p)$ with vertex set $W_{L}\cup W_{R}$ contains a $\gamma$-balanced independent set of size $(1+\epsilon-c^{\prime})\log_{b}n/\gamma$.
Proof.
Applying Theorem 3.1 to $G_{\mathrm{bip}}(n^{\theta}+O(1),p)$, it suffices to verify that
$\frac{\theta}{\gamma(1-\gamma)}\log_{b}n>(1+\epsilon-c^{\prime})\frac{\log_{b}n}{\gamma}\Leftrightarrow\theta>(1-\gamma)(1+\epsilon-c^{\prime}),$
which holds due to (25). ∎
Final Phases
For rounds $T+4\leq t\leq 2n$:
  • We randomly sample a vertex $v_{t}\in[2n]\setminus(L_{t-1}\cup R_{t-1})$. If $v_{t}\in R$, update $R_{t}=R_{t-1}\cup\{v_{t}\}$ and $L_{t}=L_{t-1}$. Otherwise, set $R_{t}=R_{t-1}$ and $L_{t}=L_{t-1}\cup\{v_{t}\}$.
  • If $v_{t}\in\mathcal{J}$, set $I_{t+1}=I_{t}\cup\{v_{t}\}$. Else, set $I_{t+1}=I_{t}$.
The algorithm above is online and constructs a $\gamma$-balanced independent set of size $(1+\epsilon)\alpha_{\rm COMP}$. We finally verify the constraint (3) on the number of queries arising in Definition 3.7.
Number of Future Queries
Note that $S_{t}=\varnothing$ except $t\in\{T+1,T+2,T+3\}$. We have:
[…]
$=|I_{T}^{(L)}|\cdot|J_{R}|+|I_{T}^{(R)}|\cdot|J_{L}|+|J_{L}|\cdot|J_{R}|$
$=\frac{2c^{\prime}(1+\epsilon-c^{\prime})(1-\gamma)}{\gamma}(\log_{b}n)^{2}+\frac{(1+\epsilon-c^{\prime})^{2}(1-\gamma)}{\gamma}(\log_{b}n)^{2}$
$=\frac{1-\gamma}{\gamma}\left((1+\epsilon)^{2}-(c^{\prime})^{2}\right)(\log_{b}n)^{2}$
$=c(\gamma,\epsilon)(\log_{b}n)^{2}.$
This completes the proof of Theorem 3.8.
@@
pages10 rangepages10 rangepages10 rangepages10 rangepages-1 rangepages9 rangepages6 rangepages1 rangepages32 rangepages14 rangepages9 rangepages56 rangepages15 rangepages10 rangepages30 rangepages5 rangepages10 rangepages26 rangepages10 rangepages33 rangepages67 rangepages12 rangepages31 rangepages51 rangepages1 rangepages8 rangepages24 rangepages30 rangepages12 rangepages38 rangepages28 rangepages169 rangepages13 rangepages10 rangepages11 rangepages14 rangepages1 rangepages1 rangepages22 rangepages35 rangepages27 rangepages36 rangepages31 rangepages-1